# Janus: Resilient and Adaptive Data Transmission for Enabling Timely and Efficient Cross-Facility Scientific Workflows

### v1

@@ Abstract.
In modern science, the growing complexity of large-scale projects has increased reliance on cross-facility workflows, where institutions share resources and expertise to accelerate discovery. These workflows often involve transferring massive data over wide-area networks. While high-speed networks like ESnet and data transfer services like Globus have improved data mobility, challenges remain. Large data volumes can strain bandwidth, TCP suffers from retransmissions due to packet loss, and traditional fault-tolerance methods like erasure coding introduce significant overhead.
@@
This paper presents Janus, a resilient and adaptive data transmission approach for cross-facility scientific workflows. Janus uses UDP, integrates erasure coding for fault tolerance, and applies error-bounded lossy compression to reduce overhead. This design enables users to balance transmission time and accuracy based on specific needs. Janus also adapts coding parameters to real-time network conditions and uses optimization models to determine ideal configurations. Experiments show that Janus significantly improves data transfer efficiency while preserving fidelity.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
@@
As scientific discovery increasingly relies on advanced technologies and instruments, vast amounts of data are being generated at an unprecedented rate. For example, the International Thermonuclear Experimental Reactor (ITER) (ITER, 2022), a project designed to demonstrate fusion energy as a large-scale, carbon-free energy source, is expected to produce 2 petabytes of data daily starting in 2035. Similarly, large-scale simulations such as the Nyx cosmology code (NYX simulation, 2019), used to simulate the evolution of matter in the universe, can generate several terabytes of data every hour on the fastest U.S. supercomputers. While much of this data is analyzed locally, it must often be transferred across wide-area networks (WANs) to facilitate global collaborations. This is particularly true for ITER, which, despite being located in France, enables scientists from around the world to conduct fusion experiments remotely. Afterward, the resulting data needs to be sent to distant researchers for further analysis, sometimes across continents. In cases requiring significant computational power, the data may need to be migrated to HPC systems, such as the U.S. Department of Energy’s supercomputers, to accelerate processing. These cross-facility workflows are not only vital for solving complex scientific problems but also contribute to the movement toward data democratization—the principle that scientific data should be accessible to a wide array of researchers, regardless of their geographic location, institutional affiliation, or available resources, fostering innovation through diverse perspectives. This underscores the critical importance of resilient, efficient, and timely data transmissions.
@@
Despite advancements in high-speed computer network infrastructures like the Energy Sciences Network (ESnet) (ESnet, 2025) and data transmission services such as Globus (Globus, 2025), several challenges remain. First, the volume of scientific data generated by experiments or simulations is immense, and transferring such large datasets across WANs can be time-consuming due to bandwidth constraints, slowing down the pace of discovery. Second, widely used transport protocols like TCP, which rely on retransmission mechanisms to handle packet loss, can significantly reduce transfer efficiency, particularly over long-distance, high-latency networks. As data sizes and transfer distances grow, retransmissions increase delays, undermining the real-time requirements of scientific workflows. Third, while techniques like erasure coding can mitigate retransmissions by offering fault tolerance, they introduce additional overhead during transmission. Furthermore, fixed configurations for fault tolerance are ill-suited to dynamic network environments, where factors like packet loss fluctuate. This static approach can lead to suboptimal performance, resulting in further inefficiencies.
@@
In this paper, we propose Janus, a resilient and adaptive data transmission approach designed to enable timely and efficient cross-facility scientific workflows. To accelerate data transmission over WANs, Janus adopts the UDP protocol instead of TCP. Since UDP lacks a built-in retransmission mechanism to recover from packet loss, Janus incorporates erasure coding to provide fault tolerance. To minimize the overhead of erasure coding, Janus further combines it with error-bounded lossy compression, a technique already in use by scientists to reduce data size without exceeding acceptable error bounds (Ainsworth et al., 2019a; Liang et al., 2018). This approach allows Janus to balance data transmission time and accuracy. Specifically, Janus can address two types of user requirements by solving different optimization problems: 1) If the user requires data with a guaranteed error bound, Janus minimizes data transmission time. 2) If the user needs data within a specified time frame, Janus minimizes the error in the received data. Additionally, as network conditions such as packet loss rate can fluctuate during transmission, applying a static level of redundancy to the data cannot achieve optimal transfer efficiency. Therefore, Janus employs adaptive algorithms that dynamically adjust erasure coding parameters based on real-time network conditions, ensuring more efficient data transmission.
@@
The contributions of this paper are as follows: 1) We develop two optimization models that determine the optimal erasure coding configurations to meet user-defined requirements for both data transmission time and accuracy. These models provide a systematic approach to balancing reliability and efficiency. 2) We design and implement two adaptive data transfer protocols capable of adjusting erasure coding parameters dynamically based on real-time variations in packet loss rates. These protocols ensure optimal performance by minimizing both transmission time and data loss. 3) We present extensive evaluations through simulations and real-world experiments, which demonstrate the effectiveness of our approach in improving the reliability and efficiency of scientific data transfers.
2. Background
2.1. Fault Tolerant Data Transmission
@@
Fault-tolerant data transmission across wide-area networks ensures reliable communication despite challenges like packet loss, corruption, or congestion. TCP, a widely used transport protocol, employs retransmission mechanisms for reliability. It tracks packets with sequence numbers and acknowledgments; if an acknowledgment is not received within a timeout, TCP retransmits the missing packet. However, retransmissions can cause significant delays and variability (Kumar, 1998; Lakshman et al., 2000; Barakat et al., 2000), making TCP unsuitable for low-latency or time-sensitive scenarios.
@@
Checksums (Stone et al., 1998; Stone and Partridge, 2000; Plummer, 1989) are another key component of fault-tolerant systems, used to detect errors in transmitted data. A checksum is computed and sent with the data packet; the receiver recalculates it to check for mismatches. If an error is found, retransmission is initiated if supported. While effective for error detection, checksums do not correct errors and rely on additional mechanisms like retransmissions for recovery.
@@
Erasure coding (Rizzo, 1997; Srouji et al., 2011; Leong and Ho, 2012) offers an alternative approach to fault tolerance in data transmission by proactively correcting errors. It divides data into fragments and adds parity fragments. Even if some fragments are lost, the original data can be reconstructed from the remaining fragments, reducing reliance on retransmissions and improving performance in high-loss networks. However, it introduces computational overhead for encoding and decoding and increases data volume due to the added parity fragments.
2.2. Scientific Data Refactoring
@@
Unlike traditional lossy compressors (Ballester-Ripoll et al., 2019; Chen et al., 2014; Lakshminarasimhan et al., 2013; Lindstrom and Isenburg, 2006; Sasaki et al., 2015; Lindstrom, 2014; Ainsworth et al., 2018, 2019b, 2019a; Liang et al., 2018), which typically adhere to a single user-defined error bound, data refactoring decomposes data into multiple components, allowing flexible reconstruction at different approximation levels. For instance, JPEG2000 (Rabbani, 2002) uses block-based wavelet transformations to support progressive data reconstruction. Similarly, ZFP (Lindstrom, 2014) provides progressive reconstruction but is limited to a fixed-rate mode. Progressive reconstruction has also been studied for cost-accuracy trade-offs in scientific visualizations (Kumar et al., 2014; Pascucci and Frank, 2001; Laney and Pascucci, 2004; Paoluzzi et al., 2004; Shamir and Pascucci, 2001), though many approaches lack guarantees on error bounds, which are critical for preserving scientific features.
@@
pMGARD (Liang et al., 2021) extends MGARD’s multilevel decomposition approach (Ainsworth et al., 2018, 2019b, 2019a) to offer an error-controlled data refactoring framework for scientific data. It leverages multigrid-based decomposition and $L^{2}$ projection to break data into multiple levels, each represented by multilevel coefficients used for reconstruction. The reconstruction error is mathematically bounded, ensuring precise error control for each coefficient. pMGARD employs bitplane encoding for multilevel coefficients, optimizing them into a hierarchical structure for progressive reconstruction. This hierarchical approach allows for efficient reconstruction with error bounds determined by the number of levels used, making pMGARD highly suitable for scientific applications.
3. Optimization in Scientific Data Transmission
@@
This section provides an overview of Janus and introduces two optimization models that enable it to determine the optimal redundancy for data transmission. All mathematical symbols used in the models are listed in Table 1.
@@
Table 1. Mathematical symbols used in our models
3.1. An Overview of Janus
@@
(1) $\varepsilon=\frac{\max(\|d-\tilde{d}\|)}{\max(\|d\|)},$
measures the error between the original data $d$ and the reconstructed data $\tilde{d}$, where $\max(|d-\tilde{d}|)$ is the maximum absolute difference and $\max(|d|)$ is the maximum value in the original data. In the worst-case scenario, when no levels are available for reconstruction, the relative error is 1.
@@
Figure 1. An illustration of our approach
During transmission, some packets may be lost, but the parity fragments allow the receiver to recover many lost data fragments without retransmission. As illustrated in Fig. 1(b), for the first level, as long as no more than 4 fragments are lost within each FTG, the data can be fully recovered. Similarly, lost fragments at levels 2 and 3 can be recovered using parity fragments. However, at level 4, if 2 fragments are lost from an FTG, recovery is not possible due to the single parity fragment. If retransmission is not enabled, only the top three levels can be used for reconstruction within error bound $\varepsilon_{3}$. The number of parity fragments at each level can be adjusted based on user requirements.
3.2. Optimization Models
@@
3.2.1. Minimizing Data Transmission Time with a Guaranteed Error Bound on the Received Data
@@
Scientists often require received data, once decompressed, to maintain an error within a specified bound to ensure that compression-induced errors don’t compromise subsequent analyses. If the user-specified error bound is $\varepsilon$ and $\varepsilon_{l}\leq\varepsilon<\varepsilon_{l-1}$, Janus must ensure the transfer of at least the first $l$ levels of the refactored data. When packet loss occurs, lost fragments can be recovered as long as the number of lost fragments does not exceed the parity fragments within each FTG. If packet loss prevents the recovery of enough levels, Janus retransmits the FTGs with unrecoverable losses using passive retransmission, as opposed to TCP’s active retransmission. This approach may require multiple retransmission rounds due to potential additional packet losses during retransmission.
@@
We develop an optimization model for Janus to minimize overall transmission time. The total size of the first $l$ levels is $S=S_{1}+S_{2}+\dots+S_{l}$, and the number of FTGs is $N=\frac{S}{(n-m)s}$. The time for initial transmission is $t+\frac{nN-1}{r}$, and the expected retransmission time depends on the number of FTGs needing retransmission in each round. When the packet loss rate is relatively low, unrecoverable data losses across FTGs can be modeled as independent events. Under this assumption, the number of FTGs with unrecoverable losses in a given round follows a binomial distribution. For instance, if the number of FTGs transmitted in the initial round of transmission is $N$, and the probability of a single FTG experiencing unrecoverable data loss is $p$, the probability that exactly $K$ out of $N$ FTGs experience such losses is given by $\binom{N}{K}p^{K}(1-p)^{N-K}$. The expected number of FTGs that will need retransmission in the first round of retransmission is $\sum\limits_{K=0}^{N}K\binom{N}{K}p^{K}(1-p)^{N-K}=Np$. In general, the expected number of FTGs requiring retransmission in the $i$-th round of retransmission can be derived as $Np^{i}$. The likelihood of requiring the $i$-th round of retransmission corresponds to the probability that at least one FTG transmitted in that round experiences unrecoverable loss, which is given by $1-(1-p)^{Np^{i}}$. Using this, the expected total time for receiving all the data fragments across the required levels can be calculated as follows:
(2) $E[T_{total}]=t+\frac{nN-1}{r}+\sum\limits_{i=1}^{\infty}[1-(1-p)^{Np^{i-1}}](t+\frac{nNp^{i}-1}{r})$
Since we always have $10<p<$ 1, when $i\to\infty$, $[1-(1-p)^{Np^{i-1}}]\to 0$ and $(t+\frac{nNp^{i}-1}{r})\to t$. Consequently, the summation term in $E[T_{total}]$ diminishes, ensuring that $E[T_{total}]$ converges as $i\to\infty$.
Now the key is to estimate $p$, the probability that the data fragments of an FTG cannot be recovered from the losses. If each FTG has $n$ fragments and the transmission rate is $r$ fragments per second, sending $n$ fragments takes $n/r$ seconds. If the network latency for transmitting a single fragment from the sender to the receiver is $t$, the total time from when the first fragment of the FTG is sent to when the last fragment is received can be expressed as $T=t+(n-1)/r$. The total number of fragments sent during $T$ is
(3) $u=\frac{T}{1/r}=\frac{t+(n-1)/r}{1/r}=rt+n-1$
Assume the packet loss rate remains constant or exhibits minimal variation over short time intervals. Under this assumption, packet losses can be effectively modeled using a Poisson distribution. Specifically, the probability of exactly $j$ out of the $u$ fragments being lost within $T$ is given by:
(4) $Pr(v=j)=\frac{(\lambda T)^{j}e^{-\lambda T}}{j!}=\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}$
These $j$ lost fragments may belong to different FTGs. If each FTG has $m$ parity fragments, given that there are $j$ ($j>m$) fragment losses during $T$, the probability that the lost data fragments in an FTG cannot be recovered is given by:
(5) $Pr(\text{unrecoverable}|v=j)=\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{u-n}{j-w}}{\binom{u}{j}}$
Therefore, the overall probability that the lost data fragments in an FTG cannot be recovered is given by:
@@
(6) $\begin{split}p=&\sum\limits_{j=m+1}^{u}Pr(\text{unrecoverable}|v=j)Pr(v=j)\\ =&\sum\limits_{j=m+1}^{u}\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{u-n}{j-w}}{\binom{u}{j}}\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}\\ =&\sum\limits_{j=m+1}^{rt+n-1}\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{rt-1}{j-w}}{\binom{rt+n-1}{j}}\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}\end{split}$
However, when the packet loss rate is high (e.g., when each FTG experiences more than one fragment loss on average) and number of parity fragments per FTG is small (e.g., $m=0$ or $m=1$), unrecoverable data losses across groups become correlated, meaning multiple FTGs within $T$ are likely to experience unrecoverable losses simultaneously. This correlation invalidates the probability $p$ calculated using Equation 6, rendering it unsuitable for use in Equation 2 to compute $E[T_{total}]$. To address this issue, we adopt an alternative approach to calculate $p$ that accounts for the correlations introduced by high packet loss rates. This refined calculation ensures that Equation 2 remains a reliable approximation for the total transmission time, even under conditions of elevated packet loss. Given that $\frac{t+(n-1)/r}{n/r}$ FTGs are in transit during $T$, the average fragment losses each of these FTGs experiences is $\lambda T/(\frac{t+(n-1)/r}{n/r})=\frac{\lambda n}{r}$. An FTG experiences unrecoverable data loss only when more than $m$ fragments within it are lost. Therefore, the probability $p$ that an FTG experiences unrecoverable data losses can be calculated as follows:
@@
(7) $\begin{split}p=&Pr(\text{more than}\ m\ \text{fragment losses in an FTG})\\ =&1-\sum\limits_{j=0}^{m}\frac{{(\frac{\lambda n}{r})}^{j}e^{-(\frac{\lambda n}{r})}}{j!}\end{split}$

### v2

@@ Abstract
In modern science, the growing complexity of large-scale scientific projects has led to an increasing reliance on cross-facility scientific workflows, where resources and expertise from multiple institutions and geographic locations are leveraged to accelerate scientific discovery. These workflows often require transmitting huge amounts of scientific data through wide-area networks. Although high-speed networks like ESnet and transfer services such as Globus have improved data mobility, several challenges remain. The sheer volume of data can overwhelm network bandwidth, widely used transport protocols such as TCP suffer from inefficiencies due to retransmissions triggered by packet loss, and existing fault-tolerance mechanisms like erasure coding introduce substantial overhead.
@@
In this paper, we propose Janus, a resilient and adaptable data transmission approach designed for cross-facility scientific workflows. Unlike traditional TCP-based methods, Janus leverages UDP, integrates erasure coding for fault tolerance, and combines it with error-bounded lossy compression to reduce overhead. This novel design allows users to balance data transmission time and accuracy, optimizing transfer performance based on specific scientific requirements. Additionally, Janus dynamically adjusts erasure coding parameters in response to real-time network conditions, ensuring efficient data transfers even in fluctuating environments. We develop optimization models for determining ideal configurations and implement adaptive data transfer protocols to enhance reliability. Through extensive simulations and real-network experiments, we demonstrate that Janus significantly improves transfer efficiency while maintaining data fidelity.
I Introduction
@@
As scientific discovery increasingly relies on advanced technologies and instruments, vast amounts of data are being generated at an unprecedented rate. For example, the International Thermonuclear Experimental Reactor (ITER) [Iter], a project designed to demonstrate fusion energy as a large-scale, carbon-free energy source, is expected to produce 2 petabytes of data daily starting in 2035. Similarly, large-scale simulations such as the Nyx cosmology code [Nyx], used to simulate the evolution of matter in the universe, can generate several terabytes of data every hour on the fastest U.S. supercomputers. While much of this data is analyzed locally, it must often be transferred across wide-area networks (WANs) to facilitate global collaborations. This is particularly true for ITER, which enables scientists from around the world to conduct fusion experiments remotely and then sends the resulting data from its location in France to the distant researchers for further analysis. This data must sometimes travel across continents. In cases requiring significant computational power, the data may need to be migrated to HPC systems, such as the U.S. Department of Energy’s supercomputers, to accelerate processing. These cross-facility workflows are not only vital for solving complex scientific problems, but also helps facilitate data democratization—the principle that scientific data should be accessible to a wide array of researchers, regardless of their geographic location, institutional affiliation, or available resources, fostering innovation through diverse perspectives. This underscores the critical importance of resilient, efficient, and timely data transmissions.
@@
Despite advancements in high-speed computer network infrastructures like the Energy Sciences Network (ESnet) [ESnet] and data transmission services such as Globus [Globus], several challenges remain. First, the immense volume of scientific data makes WAN transfers time-consuming, hindering the efficiency of scientific discovery. Second, widely used transport protocols like TCP can significantly reduce transfer efficiency, particularly over long-distance, high-latency networks due to retransmissions compensating packet loss. As data sizes and transfer distances grow, retransmissions increase delays, undermining the real-time requirements of scientific workflows. Third, while erasure coding can provide fault tolerance and reduce retransmissions, it adds transmission overhead. Fixed fault-tolerance configurations are also ill-suited to dynamic networks with fluctuating packet loss, often resulting in suboptimal performance and further inefficiencies.
@@
In this paper, we propose Janus, a resilient and adaptive data transmission approach designed to enable timely and efficient cross-facility scientific workflows. To accelerate data transmission over WANs, Janus utilizes the UDP protocol instead of TCP. Since UDP lacks a built-in retransmission mechanism to recover from packet loss, Janus incorporates erasure coding to provide fault tolerance. To minimize the overhead of erasure coding, Janus incorporates error-bounded lossy compression, a technique already in use by scientists to reduce data size without exceeding acceptable error bounds [ainsworth2019multilevel2, liang2018error]. This multi-faceted approach allows Janus to balance data transmission time and accuracy. Specifically, Janus can address two types of user requirements by solving different optimization problems: 1) If the user requires data with a guaranteed error bound, Janus minimizes data transmission time. 2) If the user needs data within a specified time frame, Janus minimizes the error in the received data. Additionally, as network conditions such as packet loss rate can fluctuate during transmission, applying a static level of redundancy to the data cannot achieve optimal transfer efficiency. Therefore, Janus employs adaptive algorithms that dynamically adjust erasure coding parameters based on real-time network conditions, ensuring more efficient data transmission.
@@
The contributions of this paper are as follows: 1) We develop two optimization models that determine the optimal erasure coding configurations to meet user-defined requirements for both data transmission time and accuracy. These models provide a systematic approach to balancing reliability and efficiency. 2) We design and implement two adaptive data transfer protocols capable of adjusting erasure coding parameters dynamically based on real-time variations in packet loss rates. These protocols ensure optimal performance by minimizing both transmission time and data loss. 3) We present extensive evaluations through simulations and real-world experiments that demonstrate the effectiveness of our approach in improving the reliability and efficiency of scientific data transfers.
II Background
II-A Fault Tolerant Data Transmission
@@
Fault-tolerant data transmission across wide-area networks ensures reliable communication despite challenges like packet loss, corruption, or congestion. TCP, a widely used transport protocol, employs retransmission mechanisms for reliability. It tracks packets with sequence numbers and acknowledgments; if an acknowledgment is not received within a timeout, TCP retransmits the missing packet. However, retransmissions can cause significant delays and variability [kumar1998comparative, lakshman2000tcp, barakat2000tcp], making TCP unsuitable for low-latency or time-sensitive scenarios.
@@
Checksums [stone1998performance, stone2000crc, plummer1989tcp] are another key component of fault-tolerant systems, used to detect errors in transmitted data. A checksum is computed and sent with the data packet; the receiver recalculates it to check for mismatches. If an error is found, retransmission is initiated if supported. While effective for error detection, checksums do not correct errors and rely on additional mechanisms like retransmissions for recovery.
@@
Erasure coding [rizzo1997effective, srouji2011rdts, leong2012erasure] offers an alternative approach to fault tolerance in data transmission by proactively correcting errors. It divides data into fragments and adds parity fragments. Even if some fragments are lost, the original data can be reconstructed from the remaining fragments, reducing reliance on retransmissions and improving performance in high-loss networks. However, erasure coding introduces computational overhead for encoding and decoding and increases data volume due to the added parity fragments.
II-B Scientific Data Refactoring
@@
Unlike traditional lossy compressors [ballester2019tthresh, chen2014numarck, lakshminarasimhan2013isabela, lindstrom2006fast, sasaki2015exploration, lindstrom2014fixed, ainsworth2018multilevel, ainsworth2019multilevel, ainsworth2019multilevel2, liang2018error] that typically adhere to a single user-defined error bound, data refactoring decomposes data into multiple components, allowing flexible reconstruction at different approximation levels. For instance, JPEG2000 [rabbani2002book] uses block-based wavelet transformations to support progressive data reconstruction. Similarly, ZFP [lindstrom2014fixed] provides progressive reconstruction but is limited to a fixed-rate mode. Progressive reconstruction has also been studied for cost-accuracy trade-offs in scientific visualizations [kumar2014efficient, pascucci2001global, laney2004progressive, paoluzzi2004progressive, shamir2001temporal], though many approaches lack guarantees on error bounds that are critical for preserving scientific features.
@@
pMGARD [liang2021error] extends MGARD’s multilevel decomposition approach [ainsworth2018multilevel, ainsworth2019multilevel, ainsworth2019multilevel2] to offer an error-controlled data refactoring framework for scientific data. It leverages multigrid-based decomposition and $L^{2}$ projection to break data into multiple levels, each represented by multilevel coefficients used for reconstruction. The reconstruction error is mathematically bounded, ensuring precise error control for each coefficient. pMGARD employs bitplane encoding for multilevel coefficients, optimizing them into a hierarchical structure for progressive reconstruction. This hierarchical approach allows for efficient reconstruction with error bounds determined by the number of levels used, making pMGARD highly suitable for scientific applications.
III Optimization in Scientific Data Transmission
@@
This section provides an overview of Janus and introduces two optimization models that enable it to determine the optimal redundancy for data transmission. All mathematical symbols used in the models are listed in Table I.
@@
TABLE I: Mathematical symbols used in our models
III-A An Overview of Janus
@@
$\varepsilon=\frac{\max(\|d-\tilde{d}\|)}{\max(\|d\|)},$ (1)
[…]
@@
Figure 1: An illustration of data transfer with Janus
During transmission, some packets may be lost, but parity fragments allow the receiver to recover many lost fragments without retransmission. As shown in Fig. 1(b), at level 1, data can be fully recovered as long as no more than 4 fragments are lost per FTG. Similarly, lost fragments at levels 2 and 3 can be recovered using parity fragments. At level 4, however, losing 2 fragments from an FTG cannot be recovered due to the single parity fragment. Without retransmission, only the top three levels can be reconstructed within the error bound $\varepsilon_{3}$. The number of parity fragments per level can be adjusted according to user requirements.
III-B Optimization Models
@@
III-B1 Minimizing Data Transmission Time with a Guaranteed Error Bound on the Received Data
@@
Scientists often require that received data maintain an error within a specified bound after decompression to ensure compression-induced errors do not compromise downstream analyses. If the user-specified bound is $\varepsilon$ and $\varepsilon_{l}\leq\varepsilon<\varepsilon_{l-1}$, Janus ensures that at least the first $l$ levels of the refactored data are transferred. Lost fragments can be recovered as long as their number does not exceed the parity fragments within each FTG. If insufficient levels are recoverable, Janus performs passive retransmission of FTGs with unrecoverable losses, which may require multiple rounds due to potential additional losses during retransmission.
@@
To minimize overall transmission time, we develop an optimization model. The total size of the first $l$ levels is $S=S_{1}+\dots+S_{l}$, and the number of FTGs is $N=\frac{S}{(n-m)s}$. The initial transmission time is $t+\frac{nN-1}{r}$, and expected retransmission time depends on the number of FTGs requiring retransmission per round. For low packet loss rates, unrecoverable losses can be modeled as independent events, where the number of FTGs with unrecoverable losses follows a binomial distribution. If $N$ FTGs are transmitted initially and the probability of unrecoverable loss in a single FTG is $p$, the expected number of FTGs requiring retransmission in the first round is $\sum_{K=0}^{N}K\binom{N}{K}p^{K}(1-p)^{N-K}=Np$. Consequently, the expected number of FTGs needing retransmission in the $i$-th round is $Np^{i}$. The probability that at least one FTG in the $i$-th round experiences unrecoverable loss is $1-(1-p)^{Np^{i}}$. Thus, the expected total transmission time is
$E[T_{total}]=t+\frac{nN-1}{r}+\sum\limits_{i=1}^{\infty}[1-(1-p)^{Np^{i-1}}](t+\frac{nNp^{i}-1}{r})$ (2)
which converges as $i\to\infty$ since $0<p<1$.
Estimating $p$, the probability that an FTG cannot be recovered, depends on the number of fragments $n$, transmission rate $r$, and network latency $t$. If T=t+(n1)/rT=t+(n-1)/r is the total time for sending an FTG, then u=rt+n1u=rt+n-1 fragments are sent during $T$. Assuming a roughly constant packet loss rate $\lambda$, fragment losses can be modeled using a Poisson distribution. Specifically, the probability of exactly $j$ out of the $u$ fragments being lost within $T$ is given by:
$Pr(v=j)=\frac{(\lambda T)^{j}e^{-\lambda T}}{j!}=\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}$ (3)
For an FTG with $m$ parity fragments, the probability of unrecoverable loss given $j>m$ lost fragments is
$Pr(\text{unrecoverable}|v=j)=\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{u-n}{j-w}}{\binom{u}{j}}$ (4)
and the overall probability is
@@
$\begin{split}p=&\sum\limits_{j=m+1}^{u}Pr(\text{unrecoverable}|v=j)Pr(v=j)\\ =&\sum\limits_{j=m+1}^{u}\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{u-n}{j-w}}{\binom{u}{j}}\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}\\ =&\sum\limits_{j=m+1}^{rt+n-1}\frac{\sum\limits_{w=m+1}^{n}\binom{n}{w}\binom{rt-1}{j-w}}{\binom{rt+n-1}{j}}\frac{{\lambda(t+\frac{n-1}{r})}^{j}e^{-\lambda(t+\frac{n-1}{r})}}{j!}\end{split}$ (5)
At high packet loss rates, unrecoverable losses across FTGs are correlated, invalidating the above calculation. We therefore we adopt an alternative approach to calculate $p$. Given that the average fragment losses each of these FTGs experiences is $\lambda T/(\frac{t+(n-1)/r}{n/r})=\frac{\lambda n}{r}$, if $\frac{\lambda n}{r}>1$ (more than one fragment loss per FTG), the probability $p$ that an FTG experiences unrecoverable data losses can be approximately estimated as:
@@
$\begin{split}p=&Pr(\text{more than}\ m\ \text{fragment losses in an FTG})\\ =&1-\sum\limits_{j=0}^{m}\frac{{(\frac{\lambda n}{r})}^{j}e^{-(\frac{\lambda n}{r})}}{j!}\end{split}$ (6)